Particles A and B carryon charge 3.0nC are at the base corners of an equilateral triangle 2.8 m on a side

How much work is required to bring a positively charged particle carrying a charge of 5.0nC from infinity to apex if A and B are fixed?
1
Expert's answer
2021-10-12T16:23:10-0400
"W=k\frac{q^2}{a}\\W=(9\cdot10^{9})\frac{(3\cdot10^{-9})^2}{2.8}=2.9\cdot10^{-8}J"